# Charge-dependent nucleon-nucleon interaction at N3LO in nuclear lattice effective field theory

### v1

@@ Abstract
The nuclear lattice effective field theory (NLEFT) is an efficient tool for solving nuclear many-body problems, which takes high-fidelity lattice chiral interactions as input and computes nuclear low-energy observables via quantum Monte Carlo techniques. In this work, we present the first next-to-next-to-next-to-leading order (N3LO) chiral forces on the lattice with the isospin-breaking effects fully taken into account. We focus on both the charge-independence breaking (CIB) and charge-symmetry breaking (CSB) effects. Specifically, we include the isospin-breaking effect from the mass difference between the charged and neutral pions in the one-pion-exchange potential (OPEP), the Coulomb force for the $pp$ interaction and the contribution of two additional charge-dependent contact operators. We also explicitly incorporate the two-pion-exchange potentials which was mostly neglected in previous NLEFT calculations. With these improvements, we are able to accurately reproduce the $np$ and $pp$ scattering phase shifts up to relative momentum $p\sim 200$ MeV as well as the deuteron properties. The construction of these charge-dependent lattice nuclear forces establishes a solid foundation for future high-precision nuclear ab initio calculations within the NLEFT framework.
@@ III Charge-independence breaking and charge-symmetry breaking
To have a clear view of this issue, in the following we list the empirical values of the scattering length $a$ and effective range $r$ for $pp$, $nn$, and $np$ scattering respectively, in the ${}^{1}S_{0}$ channel [73, 96, 97, 31],
@@ IV.1 Method to Calculate Phase Shifts
$N$ is the $n_{r}\times n_{r}$ norm matrix and $H$ is the $n_{r}\times n_{r}$ Hamiltonian defined in the $|r,LSJM\rangle$ basis, with $n_{r}$ the number of radial lattice sites.
@@ V Numerical Results
Given the lattice spacing $a$, the corresponding momentum cutoff of the lattice regularization is $\Lambda_{\mathrm{lat}}=\pi/a$. So for the small lattice spacing $a=(200~{}\mathrm{MeV})^{-1}=0.99$ fm, the momentum cutoff is $\Lambda_{\mathrm{lat}}=628$ MeV. To investigate the lattice artifacts arising from the non-zero lattice spacing, we choose several lattice spacings, i.e., $a=1.97,~{}1.64,~{}1.32$ and $0.99$ fm. Additionally, we explore the dependence on the momentum cutoff $\Lambda$ introduced for the contact terms, by using several values of $\Lambda$, i.e., $\Lambda=250,350$ and $450$ MeV.
@@ V.1 $np$ and $pp$ Scattering Phase Shifts
$C_{2,{}^{1}P_{1}}$ $0.027\pm 0.003$ $0.057\pm 0.006$ $0.14\pm 0.01$ $0.45\pm 0.04$
$C_{2,{}^{3}P_{0}}$ $-0.032\pm 0.001$ $-0.068\pm 0.003$ $-0.167\pm 0.008$ $-0.53\pm 0.02$
$C_{2,{}^{3}P_{1}}$ $0.0315\pm 0.0005$ $0.0674\pm 0.0009$ $0.165\pm 0.002$ $0.520\pm 0.007$
$C_{2,{}^{3}P_{2}}$ $-0.0436\pm 0.0007$ $-0.089\pm 0.001$ $-0.219\pm 0.003$ $-0.69\pm 0.01$
$C_{4,{}^{1}S_{0}}$ $-0.0176\pm 0.0002$ $-0.0529\pm 0.0007$ $-0.202\pm 0.002$ $-1.13\pm 0.01$
$C_{4,{}^{3}S_{1}}$ $-0.0150\pm 0.0005$ $-0.039\pm 0.002$ $-0.153\pm 0.006$ $-0.87\pm 0.03$
$C_{4,SD}$ $-0.0084\pm 0.0004$ $-0.029\pm 0.001$ $-0.111\pm 0.004$ $-0.62\pm 0.02$
$C_{4,{}^{1}P_{1}}$ $0.0000\pm 0.0007$ $0.000\pm 0.002$ $-0.001\pm 0.007$ $-0.01\pm 0.04$
$C_{4,{}^{3}P_{0}}$ $0.0035\pm 0.0004$ $0.012\pm 0.001$ $0.045\pm 0.005$ $0.26\pm 0.03$
$C_{4,{}^{3}P_{1}}$ $-0.0009\pm 0.0001$ $-0.0032\pm 0.0004$ $-0.012\pm 0.001$ $-0.067\pm 0.008$
$C_{4,{}^{3}P_{2}}$ $0.0026\pm 0.0001$ $0.0070\pm 0.0004$ $0.027\pm 0.001$ $0.149\pm 0.008$
$C_{4,PF}$ $-0.00056\pm 0.00003$ $-0.00197\pm 0.00008$ $-0.0074\pm 0.0003$ $-0.042\pm 0.002$
$C_{4,{}^{1}D_{2}}$ $-0.00169\pm 0.00004$ $-0.0051\pm 0.0001$ $-0.0193\pm 0.0004$ $-0.108\pm 0.002$
$C_{4,{}^{3}D_{1}}$ $0.0007\pm 0.0002$ $0.0020\pm 0.0007$ $0.006\pm 0.002$ $0.037\pm 0.014$
$C_{4,{}^{3}D_{2}}$ $-0.00321\pm 0.00002$ $-0.00893\pm 0.00007$ $-0.0338\pm 0.0003$ $-0.190\pm 0.001$
$C_{4,{}^{3}D_{3}}$ $-0.0003\pm 0.0001$ $-0.0028\pm 0.0003$ $-0.011\pm 0.001$ $-0.065\pm 0.008$
$C_{CD}^{pp}$ $-1.010\pm 0.001$ $-1.452\pm 0.002$ $-2.269\pm 0.003$ $-4.034\pm 0.005$
$C_{CD}^{nn}$ $-1.028\pm 0.001$ $-1.479\pm 0.002$ $-2.311\pm 0.003$ $-4.108\pm 0.005$
Figure 1: (Color online) Scattering phase shifts and mixing angles versus the relative momenta $P$ between two nucleons for the $np$ systems. The lattice spacing is set to $a=0.99$ fm, and the momentum cutoff is set to $\Lambda=350$ MeV. Notice that the TPEP is not included, which is consistent with the Monte Carlo simulations of the few- and many-body systems using NLEFT.
Figure 2: (Color online) $pp$ and $nn$ scattering phase shifts for partial wave ${}^{1}S_{0}$. Notice that the $nn$ scattering phase shifts are not used in the fits, thus the results in right panel should be taken as a lattice prediction of the $nn$ phase shifts in partial wave ${}^{1}S_{0}$.
@@ V.2 Dependence on Lattice Spacing
Figure 3: (Color online) $np$ scattering phase shifts and mixing angles versus relative momentum $P$ between neutron and proton, calculated with a lattice spacing of $a=1.97~{}{\rm fm}$. The momentum cutoff is fixed as $\Lambda=350$ MeV.
Figure 4: (Color online) $np$ phase shifts and mixing angles versus relative momentum $P$ between neutron and proton, calculated with a lattice spacing of $a=1.64~{}{\rm fm}$. The momentum cutoff is fixed as $\Lambda=350$ MeV.
Figure 5: (Color online) $np$ phase shifts and mixing angles versus the relative momentum $P$ between neutron and proton, calculated with a lattice spacing of $a=1.32~{}{\rm fm}$. The momentum cutoff is fixed as $\Lambda=350$ MeV.
@@ V.3 Deuteron Properties and $S$-wave Parameters
Table 2: Deuteron properties and $S$-wave parameters calculated with the momentum cutoff $\Lambda=350$ MeV.
@@
$E_{d}$ (MeV) $2.2606$ $2.2135$ $2.2141$ $2.2159$ 2.224575(9) [105]
$A_{s}(\mathrm{fm}^{-1/2})$ $0.8820$ $0.8729$ $0.8716$ $0.8720$ 0.8846(9) [106]
$\zeta$ $0.02220$ $0.02189$ $0.0214$ $0.0217$ 0.0256(4) [107]
$Q_{d}(\mathrm{fm}^{2})$ $0.2711$ $0.2753$ $0.2755$ $0.2748$ 0.2859(3) [108]
$r_{d}$ (fm) $1.9663$ $1.9785$ $1.9774$ $1.9733$ 1.97535(85) [109]
$a_{{}^{3}S_{1}}$ $5.408$ $5.421$ $5.421$ $5.513$ 5.424(4) [110]
$r_{{}^{3}S_{1}}$ $1.829$ $1.746$ $1.747$ $1.746$ 1.759(5) [110]
$a_{{}^{1}S_{0}}$ $-23.691$ $-23.663$ $-23.664$ $-23.664$23.748(10) [110]
$r_{{}^{1}S_{0}}$ $2.6430$ $2.6170$ $2.6173$ $2.6194$ 2.75(5) [110]
@@ V.4 Dependence on the Momentum Cutoff
As discussed in section II, we introduce a regulator $\exp(-p^{6}/(2\Lambda^{6}))$ for each single nucleon, with $\Lambda$ the momentum cutoff. To investigate the dependence on $\Lambda$, we perform calculations with several cutoffs, i.e., $\Lambda=250,350$, and $450$ MeV. In the calculation, the lattice spacing is fixed as $a=0.99$ fm. In Fig. 6, we show the plots of the $np$ scattering phase shifts with different cutoffs. One can see that except for some channels such as ${}^{1}S_{0}$, ${}^{3}P2$, and $\epsilon_{1}$, the results with different cutoffs are consistent with each other, and match the results from Nijmegan partial wave analysis for relative momentum $p<200$ MeV. For channels ${}^{1}S_{0}$, ${}^{3}P_{2}$ and $\epsilon_{1}$, the result of $\Lambda=250$ MeV has some discrepancies from others for the relative momentum $p$ around $200$ MeV. The difference between $\Lambda=350$ and 450 MeV, however, is quite small, indicting that the dependence on the momentum cutoff is negligible for $450350<\Lambda<$ 450 MeV.
@@ V.5 Contribution of Two-pion-exchange Potential
Table 3: Low-energy constants determined by N3LO fits using lattice spacings, $a=1.97~{}{\rm fm}$ , $1.64~{}{\rm fm}$, $1.32~{}{\rm fm}$ and $0.99~{}{\rm fm}$. The momentum cutoff is set to $\Lambda=350$ MeV. TPEP has been included.
@@ VI Summary and Outlook
For a lattice calculation, the application of nonzero lattice spacing induces lattice artifacts. In the present work, we investigate the lattice artifacts by performing calculations using four different lattice spacings, i.e., $a=1.97,~{}1.64,~{}1.32$ and $0.99$ fm. We find that for the coarsest lattice, $a=1.97$ fm, the $np$ phase shifts of $S$ and $P$ waves can be described pretty well for relative momentum $p<200$ MeV. However, for some higher partial waves, such as ${}^{3}D_{3}$, ${}^{3}F_{2}$, and $3G_{3}$, there are non-negligible discrepancy for $p>150$ MeV. Decreasing the lattice spacing to $1.64$ fm, $np$ phase shifts of $S$, $P$, and $D$ waves can be well described in the range $p<200$ MeV. Further decreasing the lattice spacing to $1.32$ fm, and then to $0.99$ fm, the results are stable and converge to the empirical values for all $S$, $P$, and $D$ waves for relative momentum $p<200$ MeV. This means that lattice spacings of $1.32$ and $0.99$ fm are small enough for the calculation of low-energy nuclear physics.
We regulate the contact terms nonlocally by introducing a regulator $\exp(-p^{6}/(2\Lambda^{6}))$ for each nucleon, with $\Lambda$ a momentum cutoff. We also explore the dependence on $\Lambda$ by performing calculations with different cutoffs ranging from $250$ to $450$ MeV, i.e., $\Lambda=250,350$, and $450$ MeV. We find that the dependence on the momentum cutoff is negligible for $450350<\Lambda<$ 450 MeV, whereas the results with $\Lambda=250$ MeV have some discrepancies from those of the other two when $p$ is round 200 MeV. The long-range nuclear force arises from OPEP as well as TPEP up to N3LO. It is argued that for low-energy nuclear physics TPEP is analytical and can be absorbed by redefining the LECs associated with the contact terms. We also investigate the contribution of TPEP by comparing the results with and without specific TPEP. We found that when the lattice spacing is fixed at $0.99$ fm, both results almost match each other, except for small discrepancies around $p=200$ MeV.
To draw a conclusion, we present a high-fidelity two-body nuclear force at N3LO within NLEFT, with which one can precisely generate the $np$, $pp$ and $nn$ scattering data. In the future, we will apply this interaction to the many-body simulations using NLEFT, such as the Hoyle state, as well as structures and properties of 16O.

### v2

@@ Abstract
The nuclear lattice effective field theory (NLEFT) is an efficient tool for solving nuclear many-body problems, which takes high-fidelity lattice chiral interactions as input and computes nuclear low-energy observables via quantum Monte Carlo techniques. In this work, we present the first next-to-next-to-next-to-leading order (N3LO) chiral forces on the lattice with the isospin-breaking effects fully taken into account. We focus on both the charge-independence breaking (CIB) and charge-symmetry breaking (CSB) effects. Specifically, we include the isospin-breaking effect from the mass difference between the charged and neutral pions in the one-pion-exchange potential (OPEP), the Coulomb force for the $pp$ interaction and the contribution of two additional charge-dependent contact operators. We also explicitly incorporate the two-pion-exchange potentials which was mostly neglected in previous NLEFT calculations. With these improvements, we are able to accurately reproduce the $np$ and $pp$ scattering phase shifts up to relative momentum $p\approx 200$ MeV as well as the deuteron properties. The construction of these charge-dependent lattice nuclear forces establishes a solid foundation for future high-precision nuclear ab initio calculations within the NLEFT framework.
@@ III Charge-independence breaking and charge-symmetry breaking
To have a clear view of this issue, in the following we list the empirical values of the scattering length $a$ and effective range $r$ for $pp$, nnnn, and $np$ scattering respectively, in the $^{1}S_{0}$ channel [73, 96, 97, 31],
@@ IV.1 Method to Calculate Phase Shifts
NN is the $n_{r}\times n_{r}$ norm matrix and $H$ is the $n_{r}\times n_{r}$ Hamiltonian defined in the $|r,LSJM\rangle$ basis, with $n_{r}$ the number of radial lattice sites.
@@ V Numerical Results
Given the lattice spacing $a$, the corresponding momentum cutoff of the lattice regularization is $\Lambda_{\mathrm{lat}}=\pi/a$. So for the small lattice spacing $a=(200~\mathrm{MeV})^{-1}=0.99$ fm, the momentum cutoff is $\Lambda_{\mathrm{lat}}=628$ MeV. To investigate the lattice artifacts arising from the non-zero lattice spacing, we choose several lattice spacings, i.e., a=1.97,1.64,1.32a=1.97,~1.64,~1.32 and $0.99$ fm. Additionally, we explore the dependence on the momentum cutoff $\Lambda$ introduced for the contact terms, by using several values of $\Lambda$, i.e., $\Lambda=250,350$ and $450$ MeV.
@@ V.1 $np$ and $pp$ Scattering Phase Shifts
$C_{2,{}^{1}P_{1}}$ $027\pm 0.003 0.0$$057\pm 0.006$ $0.14\pm 0.01$ $0.45\pm 0.04$
$C_{2,{}^{3}P_{0}}$ $032\pm 0.001 0.0$$068\pm 0.003$ $0.167\pm 0.008$ $0.53\pm 0.02$
$C_{2,{}^{3}P_{1}}$ $0315\pm 0.0005 0.0$$0674\pm 0.0009$ $0.165\pm 0.002$ $0.520\pm 0.007$
$C_{2,{}^{3}P_{2}}$ $0436\pm 0.0007 0.0$$089\pm 0.001$ $0.219\pm 0.003$ $0.69\pm 0.01$
$C_{4,{}^{1}S_{0}}$ $0176\pm 0.0002 0.0$$0529\pm 0.0007$ $0.202\pm 0.002$ $1.13\pm 0.01$
$C_{4,{}^{3}S_{1}}$ $0150\pm 0.0005 0.0$$039\pm 0.002$ $0.153\pm 0.006$ $0.87\pm 0.03$
$C_{4,SD}$ $0084\pm 0.0004 0.0$$029\pm 0.001$ $0.111\pm 0.004$ $0.62\pm 0.02$
$C_{4,{}^{1}P_{1}}$ $00\pm 0.0007 0.000±0.0$$\pm 0.002 0.0$$01\pm 0.007 0.0$$01\pm 0.04$
$C_{4,{}^{3}P_{0}}$ $0035\pm 0.0004 0.0$$12\pm 0.001 0.0$$045\pm 0.005$ $0.26\pm 0.03$
$C_{4,{}^{3}P_{1}}$ $0009\pm 0.0001 0.0$$032\pm 0.0004 0.0$$12\pm 0.001 0.0$$067\pm 0.008$
$C_{4,{}^{3}P_{2}}$ $0026\pm 0.0001 0.0$$070\pm 0.0004 0.0$$027\pm 0.001$ $0.149\pm 0.008$
$C_{4,PF}$ $00056\pm 0.00003 0.0$$0197\pm 0.00008 0.0$$074\pm 0.0003 0.0$$042\pm 0.002$
$C_{4,{}^{1}D_{2}}$ $00169\pm 0.00004 0.0$$051\pm 0.0001 0.0$$0193\pm 0.0004$ $0.108\pm 0.002$
$C_{4,{}^{3}D_{1}}$ $0007\pm 0.0002 0.0$$020\pm 0.0007 0.0$$06\pm 0.002 0.0$$037\pm 0.014$
$C_{4,{}^{3}D_{2}}$ $00321\pm 0.00002 0.0$$0893\pm 0.00007 0.0$$0338\pm 0.0003$ $0.190\pm 0.001$
$C_{4,{}^{3}D_{3}}$ $0003\pm 0.0001 0.0$$028\pm 0.0003 0.0$$11\pm 0.001 0.0$$065\pm 0.008$
$C_{CD}^{pp}$ $1.010\pm 0.001$ $1.452\pm 0.002$ $2.269\pm 0.003$ $4.034\pm 0.005$
$C_{CD}^{nn}$ $1.028\pm 0.001$ $1.479\pm 0.002$ $2.311\pm 0.003$ $4.108\pm 0.005$
Figure 1: (Color online) Scattering phase shifts and mixing angles versus the relative momenta $P$ between two nucleons for the npnp systems. The lattice spacing is set to $a=0.99$ fm, and the momentum cutoff is set to $\Lambda=350$ MeV. Notice that the TPEP is not included, which is consistent with the Monte Carlo simulations of the few- and many-body systems using NLEFT.
Figure 2: (Color online) pppp and $nn$ scattering phase shifts for partial wave $^{1}S_{0}$. Notice that the $nn$ scattering phase shifts are not used in the fits, thus the results in right panel should be taken as a lattice prediction of the $nn$ phase shifts in partial wave $^{1}S_{0}$.
@@ V.2 Dependence on Lattice Spacing
Figure 3: (Color online) npnp scattering phase shifts and mixing angles versus relative momentum $P$ between neutron and proton, calculated with a lattice spacing of $a=1.97~{\rm fm}$. The momentum cutoff is fixed as $\Lambda=350$ MeV.
Figure 4: (Color online) npnp phase shifts and mixing angles versus relative momentum $P$ between neutron and proton, calculated with a lattice spacing of $a=1.64~{\rm fm}$. The momentum cutoff is fixed as $\Lambda=350$ MeV.
Figure 5: (Color online) npnp phase shifts and mixing angles versus the relative momentum $P$ between neutron and proton, calculated with a lattice spacing of $a=1.32~{\rm fm}$. The momentum cutoff is fixed as $\Lambda=350$ MeV.
@@ V.3 Deuteron Properties and $S$-wave Parameters
Table 2: Deuteron properties and $S$-wave parameters calculated with the momentum cutoff $\Lambda=350$ MeV. Notice that the deuteron binding energy is not used as input in the fits.
@@
$E_{d}$ (MeV) $26062\pm 0.00004 2.2$$1354\pm 0.00004 2.2$$1407\pm 0.00003 2.2$$1595\pm 0.00002 2.2$24575(9) [105]
$A_{s}(\mathrm{fm}^{-1/2})$ $0.882\pm 0.001$ $0.873\pm 0.001$ $0.872\pm 0.001$ $0.872\pm 0.001$ 0.8846(9) [106]
$\zeta$ $02220\pm 0.00008 0.0$$2189\pm 0.00009 0.0$$2138\pm 0.00009 0.0$$2170\pm 0.00009 0.0$256(4) [107]
$Q_{d}(\mathrm{fm}^{2})$ $0.271\pm 0.002$ $0.275\pm 0.002$ $0.276\pm 0.002$ $0.275\pm 0.002$ 0.2859(3) [108]
$r_{d}$ (fm) $1.9663\pm 0.0009$ $1.9785\pm 0.0009$ $1.9774\pm 0.0009$ $1.9733\pm 0.0009$ 1.97535(85) [109]
$a_{{}^{3}S_{1}}$ $5.408\pm 0.003$ $5.420\pm 0.003$ $5.421\pm 0.003$ $5.421\pm 0.003$ 5.424(4) [110]
$r_{{}^{3}S_{1}}$ $1.829\pm 0.008$ $1.746\pm 0.003$ $1.747\pm 0.002$ $1.746\pm 0.002$ 1.759(5) [110]
$a_{{}^{1}S_{0}}$ $23.69\pm 0.06$ $23.71\pm 0.06$ $23.71\pm 0.06$ $23.71\pm 0.06 -$23.748(10) [110]
$r_{{}^{1}S_{0}}$ $2.643\pm 0.002$ $2.646\pm 0.002$ $2.649\pm 0.002$ $2.650\pm 0.002$ 2.75(5) [110]
@@ V.4 Dependence on the Momentum Cutoff
As discussed in section II, we introduce a regulator $\exp(-p^{6}/(2\Lambda^{6}))$ for each single nucleon, with $\Lambda$ the momentum cutoff. To investigate the dependence on $\Lambda$, we perform calculations with several cutoffs, i.e., $\Lambda=250,350$, and $450$ MeV. In the calculation, the lattice spacing is fixed as $a=0.99$ fm. In Fig. 6, we show the plots of the $np$ scattering phase shifts with different cutoffs. One can see that except for some channels such as $^{1}S_{0}$, $^{3}P2$, and $\epsilon_{1}$, the results with different cutoffs are consistent with each other, and match the results from Nijmegan partial wave analysis for relative momentum $p<200$ MeV. For channels $^{1}S_{0}$, $^{3}P_{2}$ and $\epsilon_{1}$, the result of $\Lambda=250$ MeV has some discrepancies from others for the relative momentum $p$ around $200$ MeV. The difference between $\Lambda=350$ and 450 MeV, however, is quite small, indicting that the dependence on the momentum cutoff is negligible for $350<\Lambda<450$ MeV.
@@ V.5 Contribution of Two-pion-exchange Potential
Table 3: Low-energy constants determined by N3LO fits using a lattice spacing, $a=0.99~{\rm fm}$. The momentum cutoff is set to $\Lambda=350$ MeV. TPEP has been included.
@@ VI Summary and Outlook
For a lattice calculation, the application of nonzero lattice spacing induces lattice artifacts. In the present work, we investigate the lattice artifacts by performing calculations using four different lattice spacings, i.e., a=1.97,1.64,1.32a=1.97,~1.64,~1.32 and $0.99$ fm. We find that for the coarsest lattice, $a=1.97$ fm, the $np$ phase shifts of $S$ and $P$ waves can be described pretty well for relative momentum $p<200$ MeV. However, for some higher partial waves, such as $^{3}D_{3}$, $^{3}F_{2}$, and $3G_{3}$, there are non-negligible discrepancy for $p>150$ MeV. Decreasing the lattice spacing to $1.64$ fm, $np$ phase shifts of $S$, $P$, and $D$ waves can be well described in the range $p<200$ MeV. Further decreasing the lattice spacing to $1.32$ fm, and then to $0.99$ fm, the results are stable and converge to the empirical values for all $S$, $P$, and $D$ waves for relative momentum $p<200$ MeV. This means that lattice spacings of $1.32$ and $0.99$ fm are small enough for the calculation of low-energy nuclear physics.
We regulate the contact terms nonlocally by introducing a regulator $\exp(-p^{6}/(2\Lambda^{6}))$ for each nucleon, with $\Lambda$ a momentum cutoff. We also explore the dependence on $\Lambda$ by performing calculations with different cutoffs ranging from $250$ to $450$ MeV, i.e., $\Lambda=250,350$, and $450$ MeV. We find that the dependence on the momentum cutoff is negligible for $350<\Lambda<450$ MeV, whereas the results with $\Lambda=250$ MeV have some discrepancies from those of the other two when $p$ is round 200 MeV. The long-range nuclear force arises from OPEP as well as TPEP up to N3LO. It is argued that for low-energy nuclear physics TPEP is analytical and can be absorbed by redefining the LECs associated with the contact terms. We also investigate the contribution of TPEP by comparing the results with and without specific TPEP. We found that when the lattice spacing is fixed at $0.99$ fm, both results almost match each other, except for small discrepancies around $p=200$ MeV.
To draw a conclusion, we present a high-fidelity two-body nuclear force at N3LO within NLEFT, with which one can precisely generate the $np$, pppp and $nn$ scattering data. In the future, we will apply this interaction to the many-body simulations using NLEFT, such as the Hoyle state, as well as structures and properties of 16O.